### v1

# Beyond Syntax: Action Semantics Learning for App Agents

Abstract
The advent of Large Language Models (LLMs) enables the rise of App agents that interpret user intent and operate smartphone Apps through actions such as clicking and scrolling. While prompt-based solutions with closed LLM APIs show promising ability, they incur heavy compute costs and external API dependency. Fine-tuning smaller open-source LLMs solves these limitations. However, current fine-tuning methods use a syntax learning paradigm that forces agents to reproduce exactly the ground truth action strings, leading to out-of-distribution (OOD) vulnerability. To fill this gap, we propose Action Semantics Learning (ASL), a novel learning framework, where the learning objective is capturing the semantics of the ground truth actions. Specifically, inspired by the programming language theory, we define the action semantics for App agents as the state transition induced by the action in the user interface. With this insight, ASL employs a novel SEmantic Estimator (SEE) to compute a semantic reward to train the App agents in generating actions aligned with the semantics of ground truth actions, even when the syntactic forms differ. To support the effectiveness of ASL, we theoretically demonstrate the superior robustness of ASL for the OOD problem compared with the existing syntax learning paradigm. Extensive experiments on offline and online smartphone App operation benchmarks show that ASL significantly improves the accuracy and generalisation of App agents over existing methods.
1 Introduction
The rapid development of Large Language Models (LLMs)  [^1] In this work, LLMs refer to foundation models that process multiple input modalities (e.g., visual or multimodal) and generate textual outputs. has led to the emergence of AI agents across a wide range of domains li2023camel ; goutora ; ramos2025review ; hu2025selfevolving . Among these, AI agents for smartphone App operation, commonly referred to as App agents, are gaining significant attention for leveraging LLMs to enable fully automated App navigation rawles2024androidinthewild ; rawles2024androidworld ; mobileagentv2 ; wang2025distrl ; chen2025spa ; christianos2025lightweight . These agents interpret user-provided goals in natural language and execute tasks such as scheduling, messaging, and online purchasing by interacting with the user interfaces (UIs) via human-like actions, such as tapping, scrolling, or entering text. The diversity of smartphone operating systems and UI designs, coupled with predefined action sets for specific systems, poses a significant challenge in developing generalisable App agents.
Recent advances in App agents can be broadly grouped into two approaches. One prominent approach leverages closed LLMs APIs, such as GPT and Gemini, through prompt engineering techniques to comprehend user intents and device states, subsequently generating suitable actions rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . While these App agents exhibit robust reasoning and generalisation capabilities, their practical deployment is often hindered by challenges such as high inference latency, substantial computational resource requirements, and frequent dependence on third-party APIs, which may escalate operational costs and introduce instability. Alternatively, fine-tuning smaller LLMs with offline datasets or interactive online trajectories has emerged as a viable strategy, achieving low-latency inference conducive to real-world applications wang2025distrl ; christianos2025lightweight ; bai2024digirl . However, current methods typically follow a syntax learning paradigm that forces agents to reproduce exactly the ground truth action string for each UI within the training data and ignores the underlying action semantics. This leads to an out-of-distribution (OOD) vulnerability: even the UIs slightly differ from the samples within the training data can provoke a dramatic performance drop. An agent trained to follow a specific interface path may fail when the UI changes or alternative options are presented. For example, an agent trained to delete text by generating "Tap the Delete button" will fail if the Delete button is unavailable, whilst the "Cut" button is still visible. See Sec. 3.1 for further discussion of this OOD vulnerability.
Figure 1: Our action semantics learning framework.
To go beyond syntax and capture action semantics, we first draw inspiration from the denotational semantics in programming language theory, which defines program semantics by its effect on system state stoy1981denotational , to define the action semantics for App agents as the state transition induced by the action in the UI environment. Building on this insight, we propose Action Semantics Learning (ASL), a novel learning framework that shifts the learning objective of App agents from exactly reproducing the ground truth action strings to reproducing their semantics. Specifically, during training, the actions generated by the App agent are rewarded by a SEmantic Estimator (SEE) that scores how closely the semantics of the generated action match that of the ground truth action. With this semantic reward, ASL allocates credit to all actions yielding the same UI state transition as the ground truth action, teaching the agent to prioritise action semantics over action strings. In Sec. 3.4, we theoretically demonstrate that, compared with the syntax learning paradigm, ASL is more robust towards the OOD problem. See Fig. 1 for a visualisation of the proposed ASL framework.
The semantic estimator SEE is the core component of our learning framework. Specifically, SEE leverages an LLM-driven world model to predict the impact of an action on the current UI state and uses BERT to measure the semantic similarity between the effects of the predicted action and the ground truth action. Our design is inspired by recent world-model–empowered web agents web-world , which leverage predicted next observations as offline rewards for prompt engineering. In contrast, we are the first to incorporate these world-model predictions directly into the app agent training stage, generating a semantic reward without requiring any extra computational resources for deployment.
Our main contributions are summarised as follows:
$\bullet$ We introduce ASL, a novel learning framework based on action semantics. In this framework, we formulate the App-agent training as the semantics-level learning: an action is correct if it produces the target UI state transition, regardless of its exact syntactic expression. This learning objective contributes to improving the generalisability of App agents.
$\bullet$ We propose SEE, a novel training-free semantic estimator that generates reward signals based on semantic alignment, enabling the fine-tuning of lightweight models to learn action semantics without additional GPU memory requirements for reward model training.
$\bullet$ We conduct extensive experiments on both offline and online smartphone operation benchmarks. Empirical results demonstrate that the proposed ASL framework achieves superior accuracy and generalisation compared to existing approaches.
2 Related Works
2.1 Prompt-driven App Agents
Prompt-driven App agents mark a substantial advancement in human–device interaction, enabling natural language commands to be translated into precise UI actions on smartphones, tablets, and other mobile devices. Recent research has progressively refined these systems, with pioneering implementations like AppAgent yang2023appagent demonstrating that large pretrained models can effectively map screenshots or accessibility trees to human-like actions. Building on these foundations, more recent frameworks have introduced advanced prompting strategies incorporating planning, self-reflection, and tool use to tackle increasingly complex tasks rawles2024androidworld ; mobileagentv2 . Despite their strong performance, current prompt-driven App agents remain heavily dependent on closed LLM APIs such as GPT and Gemini rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . This dependence introduces several limitations, including high inference latency, substantial computational costs, and reliance on third-party services that raise concerns around data privacy, service reliability, and vendor lock-in. These challenges hinder real-world deployment and highlight the need for open, efficient, and controllable alternatives.
2.2 Fine-tuned App Agents
Recent research has increasingly focused on fine-tuning smaller models tailored for mobile environments. LiMAC (christianos2025lightweight, ) introduces a compact action transformer that predicts UI-grounded actions paired with a fine-tuned vision–language model for textual reasoning, while InfiGUIAgent (infiguiagent, ) adopts a two-stage approach that first learns UI semantics from screenshots and then generates actions conditioned on user instructions. Despite architectural differences, these methods follow a syntax learning paradigm that forces agents to reproduce the exact ground truth action string. This paradigm can lead to OOD vulnerability: even slight UI differences from the training data samples can provoke dramatic performance degradation. Complementing these offline approaches, several works explore online optimisation to improve generalisability and robustness to OOD data. DigiRL (bai2024digirl, ) proposes a reinforcement learning framework that simulates App control scenarios and fine-tunes policies from an offline-supervised starting point. DistRL (wang2025distrl, ) enhances this approach with asynchronous online training, boosting sample efficiency. However, these online approaches often require extensive interaction with real mobile device environments during training, significantly increasing computational and operational costs. For instance, DigiRL’s framework involves building a scalable and parallelisable Android learning environment equipped with a vision-language model-based evaluator, necessitating substantial resources for real-device emulators. Similarly, DistRL employs centralised training with decentralised data acquisition, requiring multiple device interactions to collect diverse trajectories. In our work, we aim to enhance model generalisability without requiring interaction with real device environments during training.
2.3 World Model boosted App Agents
World models simulate future states of the environment based on previous states and given actions lecun2022path ; ding2024understanding . In domains such as robotics, world models yang2023learning ; zhen20243d ; zhou2024robodreamer ; ajay2023compositional can predict future visual states, such as images or videos across diverse environments. These predictive capabilities are critical for enabling robots to understand their surroundings, make informed decisions, and accurately execute tasks. Building on these advances, researchers have extended world modelling to web environments, where models web-world ; isyour ; liu2023picture generate textual predictions of a website’s future state following a user interaction, enabling the development of more accurate web agents. However, existing applications of world models in this setting are typically used for offline feedback, which incurs additional inference overhead and is suboptimal for practical deployment. In our work, we instead leverage world models as a reward signal to fine-tune the agent by computing the semantic similarity between predicted and actual actions, requiring no additional inference time or computational resources during deployment.
3 Methodology
In this section, we begin with the problem formulation, which introduces the smartphone app operation task and summarises the key challenge for current fine-tuning methods. Then, we propose a novel learning framework named Action Semantic Learning (ASL) for App agents. Furthermore, we elaborate on the semantic estimator, the key component within the ASL framework. Finally, we conduct a theoretical analysis to support the effectiveness of the ASL framework.
3.1 Problem Formulation
Smartphone operation. We formulate smartphone operation as a Partially Observable Markov Decision Process defined by $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O},\gamma)$, where $\mathcal{S}$ represents smartphone system states, $\mathcal{A}$ denotes actions for operating UIs, $\mathcal{T}$ is the state transition function, $\mathcal{R}$ is the utility function, $\mathcal{O}$ maps states to UI observations, and $\gamma$ is the discount factor. At each timestep $t$, the App agent, which is parameterised by $\theta$ and denoted as $\pi_{\theta}$, receives an observation $o_{t}$ of the current UI state, the user goal $g$ in natural language, and action history $h_{t}$. The App agent produces an action according to $a^{\theta}_{t}=\pi_{\theta}(g,h_{t},o_{t})$, where the observation $o_{t}$ may include a subset of available screenshots and textual information like UI element trees android_control . Let $\mathcal{D}_{train}=\{(g_{i},h_{i},o_{i},a_{i}^{*})\}_{i=1}^{N}$ denote a training dataset with $N$ step samples, where $a_{i}^{*}$ is the ground truth action for sample $i$. Our objective is to fine-tune an App agent to maximise the success rate of completing the user goal $g$, while also considering generalisation and robustness. For simplicity, we assume $\gamma=1$. The utility function $\mathcal{R}$ returns 1 when the user goal can be completed within the required maximum action steps, and 0 otherwise.
OOD vulnerability for syntax learning. Current approaches for training App agents typically employ a syntax learning paradigm, where models learn to reproduce exactly the ground truth action string rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . Mathematically, their learning objective can be formulated as:
$\arg\min_{\theta}\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}[-\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})],$ (1)
where $\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\log p_{\theta}(a_{i}^{*}\mid g_{i},h_{i},o_{i})$ is the supervised fine-tuning (SFT) loss function for the data sample $i$ and $p_{\theta}(a_{i}^{*}|g_{i},h_{i},o_{i})$ is the likelihood of the ground truth action under the App agent policy $\pi_{\theta}$ for the given inputs. The key limitation of this paradigm is that it rewards the model only when the output exactly matches the ground truth action, while penalising all other outputs, even those that share identical semantics with the ground truth action. Consequently, the model optimises for token-level cross-entropy rather than understanding action semantics, leading to severe out-of-distribution (OOD) vulnerability: even slight UI differences from the training samples can cause performance drops. For example, Table 3 in Sec. 4 shows that on the AndroidControl benchmark android_control , the accuracy of the App agent fine-tuned with Llama-3 plunges from 65.5% on the in-distribution dataset (IDD) to just 57.1% on the App-Unseen OOD set. Mathematically, we formulate this OOD vulnerability as Theorem 3.1 with the proof in the Appendix.
Theorem 3.1.
Let $\pi^{\mathrm{SFT}}_{\theta}$ be an App agent trained under the syntax-learning objective of Eq. (1); for an $n$-step task with ground truth actions $\{a_{1}^{*},\dots,a_{n}^{*}\}$, define $P(\text{success})=\Pr[\text{the agent completes all $n$ steps correctly}]$ and, for any semantically preserving but syntactically non-trivial perturbation $\delta$ satisfying $\delta(a_{i}^{*})\neq a_{i}^{*}$ for at least one $i$, define $P(\text{success}\mid\delta)=\Pr[\text{the agent completes all $n$ steps correctly when each }a_{i}^{*}\text{ is replaced by }\delta(a_{i}^{*})]$; letting $\Delta(\delta)=P(\text{success})-P(\text{success}\mid\delta)$, we have $\Delta(\delta)>0$.
3.2 Action Semantics Learning
To solve this OOD vulnerability, we introduce Action Semantics Learning (ASL), where the App agents aim to learn the semantics of ground truth actions, as displayed in Fig. 1.
In this subsection, we begin by formalising the concept of action semantics for App agents. Building on this concept, we derive a semantics-aware loss function that guides the agent to learn the semantics of ground truth actions. Finally, we introduce the semantics-aware fine-tuning pipeline of our ASL framework, which leverages the semantics-aware loss function to train App agents.
Figure 2: Examples of semantically equivalent actions: (a) and (b) lead to the same GUI state, while (c) results in a different outcome.
Action semantics. Inspired by the denotational semantics in programming language theory, which defines program semantics by its effect on system state stoy1981denotational , we mathematically define the action semantics for App agents as the following:
Definition 3.1.
(Action Semantic). Let $\mathcal{S}$ be the set of concrete smartphone states and $\mathcal{A}$ the set of finite actions. Moreover, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ denotes the state transition function mapping a state-action pair to a probability distribution over next states. $S(a)\coloneqq\mathcal{T}(s,a)$.
Intuitively, the semantics of an action generated by the App agents is defined as the state transition induced by the action in the UI environment. This definition is supported by the observation that, in interactive systems such as mobile UIs, the meaning of an action is not fully captured by the action string, but rather by the transformation it causes in the environment. For instance, tapping different buttons may lead to the same screen, while actions with similar strings may have vastly different effects depending on the context, as shown in Fig. 2. Therefore, a semantic understanding of action must go beyond symbolic identifiers and focus on the consequences of the action.
Semantics-aware loss function. With the Definition 3.1, for a ground truth action $a_{i}^{\star}$, we define its semantic equivalence set as $\mathcal{A}_{a_{i}^{\star}}^{eq}=\{a\in\mathcal{A}~{}|~{}S(a)=S(a_{i}^{\star})\}$. Instead of forcing the App agent to reproduce the exact $a_{i}^{\star}$, we encourage the model to assign high probability to all actions within $\mathcal{A}_{a_{i}^{\star}}^{eq}$. In real-world scenarios, $\mathcal{A}_{a^{\star}_{i}}^{eq}$ is typically unavailable, as identifying all semantically equivalent actions requires exhaustive human annotation and environment interaction. To address this, we formulate the learning objective in terms of a semantic reward:
$J(\theta)=\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big{[}\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\big{]}\Big{]},$ (2)
where $r(\cdot)$ is a reward model that quantifies the probability that action $a$ is semantically equivalent to $a_{i}^{\star}$ given the input $(g_{i},h_{i},o_{i})$. In practice, we employ a training-free reward model. We elaborate on the design of this reward model in Sec. 3.3.
To optimise the objective function in Eq. (2), we leverage the REINFORCE algorithm (williams1992simple, ). This yields the following gradient estimator:
$\nabla_{\theta}J(\theta)\;=\;\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big{[}\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\,\nabla_{\theta}\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big{]}\Big{]},$
where $p_{\theta}(a|g_{i},h_{i},o_{i})$ is the likelihood of an action under the App agent $\pi_{\theta}$ for the given inputs. Then, we design the semantics-aware loss function for a data sample $i$ as:
$\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\,\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big{]}.$ (3)
In Sec. 3.4, we show that minimising the semantics-aware loss function can lead to an App agent that is more robust to the OOD vulnerability than the one fine-tuned with the syntax learning paradigm.
Semantics-aware fine-tuning. Optimising the semantics-aware loss function in Eq. (3) enables App agents to learn action semantics by rewarding semantically equivalent actions. However, computing this loss requires sampling actions from the agent and scoring these actions using a reward model, which can lead to prohibitively high training costs.
To maintain efficiency, inspired by previous works in the literature nakano2021webgpt ; ouyang2022training ; rafailov2023direct , we limit the number of sampled actions as one per data sample. However, sampling only a single action can lead to vanishing gradients when the sampled action is not semantically equivalent to the ground truth, hindering effective learning and potentially leading to sub-optimal performance. To address this, we propose a combined loss function that integrates our semantics-aware loss function with the SFT loss function. Concretely, for each data sample, the loss function is formulated as follows:
$\ell^{ASL}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})+\lambda_{i}\,\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*}),$ (4)
where $\lambda_{i}=\big{(}1-(1-\alpha_{i})\,r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\big{)}$, and $\alpha_{i}=\exp\big{(}-\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})\big{)}$. We use $\alpha_{i}$ to quantify the syntax dissimilarity between the action generated by the App agent and the ground truth action. The SFT loss function value used to compute $\alpha_{i}$ is detached from the gradient computational graph, ensuring it only modulates the SFT term without creating unintended gradient interactions.
Intuitively, the combined loss function operates as follows: (1) when the sampled action is semantically correct but syntactically different, learning is primarily guided by the semantic-aware loss; (2) when the action matches both semantics and syntax, both loss components contribute jointly; and (3) when the action matches neither, the SFT loss dominates, ensuring stable gradient flow even in the absence of semantic reward. This design effectively mitigates gradient vanishing issues while preserving training efficiency, as it requires sampling only a single action from the App agent per training instance. We provide a more detailed discussion for this intuition in the Appendix.
In the next section, we propose a novel SEmantic Estimator (SEE) for App agents, which serves as the reward model in our ASL framework.
3.3 Semantic Estimator
During offline training, the direct interaction with a real smartphone UI environment is often impractical. Therefore, it is challenging to get accurate semantic rewards, namely the probability that an agent-generated action is semantically equivalent to the ground truth action. To address this limitation, we introduce a SEmantic Estimator (SEE) that consists of two key components: (1) a world model that simulates the smartphone UI state transitions, and (2) a semantic calculator that estimates the semantic reward by quantifying the similarity between the transitions.
World model. Inspired by recent advances in world-model–based web agents web-world , which leverage predicted future observations as rewards for prompt engineering, SEE utilises a world model to mimic the state transition function $\mathcal{T}(s,a)$ mentioned in Definition 3.1. Mathematically, the workflow of this world model can be formulated as:
$\Delta\hat{o}_{t}=\Phi(o_{t},a_{t}),$ (5)
where $o_{t}$ is the textual description of the UI state $s_{t}$, $a_{t}$ is the textual description of an action, and $\Delta\hat{o}_{t}$ is the description of the impact of the action $a_{t}$ on the UI state $s_{t}$. In practice, we implement this world model based on the closed LLM API by leveraging the LLM’s strong capabilities in language understanding and world knowledge to simulate how the environment would evolve in response to the action. The prompt of our world model can be found in the Appendix.
Semantic calculator. With the state transitions generated by the world model, SEE employs a semantic calculator to generate the semantic reward for our ASL framework-generated transitions in two steps. Firstly, the semantic calculator uses BERT devlin2019bert to project the generated textual description for state transitions into a feature space. This process can be formulated as:
$f^{gt}_{t}=\text{BERT}(\Delta\hat{o}^{gt}_{t}),~{}~{}~{}~{}f^{pre}_{t}=\text{BERT}(\Delta\hat{o}^{pre}_{t}),$ (6)
where $\Delta\hat{o}^{gt}_{t}$ and $\Delta\hat{o}^{pre}_{t}$ denote the textual descriptions of the ground truth and predicted state transitions at timestep $t$, respectively, and $f^{gt}_{t},f^{pre}_{t}\in\mathbb{R}^{d}$ are their corresponding feature representations. Then, SEE computes the cosine similarity between the feature vectors to measure semantic alignment:
$cos_{t}=\frac{\langle f^{gt}_{t},f^{pre}_{t}\rangle}{|f^{gt}_{t}|\cdot|f^{pre}_{t}|}.$ (7)
This similarity score is used in our ASL framework as the semantic reward. A higher similarity score indicates a higher probability that a semantic equivalence exists between the action generated by the App agent and the ground truth action, providing a reward signal that guides the training of the agent towards semantically meaningful actions, even when the syntax may differ.
3.4 Theoretical Analysis
In this section, we demonstrate that enabling the App agent to learn action semantics by minimising the semantic loss function defined in Eq. (3) leads to an agent that is more robust to the OOD problems compared to an agent that purely optimises the learning objective formulated in Eq. (1). Mathematically, we present Theorem 3.2 with the proof in the Appendix.
Theorem 3.2.
Let $\pi^{\mathrm{S}}_{\theta}$ be an App agent trained by minimising the semantics-aware loss function defined in Eq. (3), and let $\pi^{\mathrm{SFT}}_{\theta}$ be an App agent trained under the syntax-learning objective defined in Eq. (1). For an $n$-step task with ground truth actions $\{a_{1}^{*},\dots,a_{n}^{*}\}$, consider any semantically preserving but syntactically non-trivial perturbation $\delta$ such that $\delta(a_{i}^{*})\neq a_{i}^{*}$ for at least one $i$. Define $P^{\mathrm{S}}(\text{success}\mid\delta)$ as the success rate of $\pi^{\mathrm{S}}_{\theta}$ and $P^{\mathrm{SFT}}(\text{success}\mid\delta)$ as the success rate of $\pi^{\mathrm{SFT}}_{\theta}$ under the perturbation $\delta$. Then, we have $P^{\mathrm{S}}(\text{success}\mid\delta)-P^{\mathrm{SFT}}(\text{success}\mid\delta)>0$.
This theorem highlights that, by focusing on the semantic equivalence of actions rather than exact syntactic matches, the App agent trained with Eq. (3) exhibits higher success rates than an agent trained solely on the syntax-learning objective when encountering the OOD data. We further provide experimental support for our learning framework in Sec. 4.
4 Experiments
In this section, we first introduce our experimental setup. We then demonstrate the practical value of the proposed ASL framework on two online smartphone operation Apps benchmarks. To evaluate the robustness of the agent fine-tuned by our ASL framework, we use an offline smartphone operation benchmark that includes a clear separation between in-distribution and OOD data. Finally, we conduct an ablation study to study the contribution of components within the ASL framework.
4.1 Experimental Setup
Datasets. We conducted our experiments on three benchmark datasets: AndroidLab xu2024androidlab , AndroidWorld rawles2024androidworld , and AndroidControl android_control . AndroidWorld includes a high-fidelity simulator that emulates realistic Android device behaviour, enabling interactions with 20 real-world applications across 116 user-defined tasks. To determine whether a task is successfully completed, each task is paired with manually authored rules that provide precise success criteria. AndroidLab includes 138 predefined tasks across nine Android applications built on virtual devices. Each task is divided into multiple required page states as sub-goals, with UI tree structure matching verifying correct traversal. The AndroidLab environment ensures reproducibility and eliminates external network or time dependencies. AndroidControl comprises 15,283 human demonstrations spanning 14,548 unique tasks across 833 Android applications. contains 15,283 human demonstrations covering 14,548 unique tasks across 833 Android applications. This dataset is divided into a training set, an in-distribution (IDD) test set, a task-level OOD test set, a category-level OOD test set, and an app-level OOD test set. Each sample consists of a user goal, an observation, and a ground-truth action.
Metrics. For the online datasets AndroidWorld and AndroidLab, where the App agent is required to complete user-specified goals within an Android emulator, we evaluate performance using the task-level success rate (TSR), which is defined as the number of successfully completed tasks divided by the total number of tasks. A task is considered successful if the agent completes the goal within the dataset’s predefined maximum number of steps, and a failure otherwise. For the offline dataset AndroidControl, where the agent uses a user goal, an observation, and an action history as inputs to generate an action matching a predefined target, we use the step-level success rate (SSR), which is defined as the number of correctly predicted steps divided by the total number of steps. A step is successful if the generated action exactly matches the ground truth action; otherwise, it is a failure.
Implementation details. For the semantic evaluator, we use Gemini-Flash-2.0 gemini2024 as the LLM to predict the change that an action will bring to the UI state. The initial learning rate was set to 5e-5, with a linear decay schedule. Batch size was set to 1 and trained for 3 epochs. Mixed precision training (fp16) was enabled to reduce memory consumption and accelerate training. The training dataset includes 2K human-annotated step-wise samples from apps in the AndroidWorld emulator environment and 70K samples from the AndroidControl benchmark. The 2K human-generated samples are curated to ensure no overlap with the OOD test set in AndroidControl. We adopt a modified version of the prompt used in T3A, a leading prompt-driven App agent that operates entirely on textual inputs. While T3A uses two separate prompts, one for the action agent and another for the action summary agent that generates action summaries for use as history, our design uses a single prompt that instructs the agent to output both the action for app interaction and a concise textual description of the action, which is reused as the action history. In contrast to T3A, which requires the agent to provide a rationale for each action, our agent only generates the action and its description. This results in a fully text-based agent that is computationally efficient and does not rely on additional modules for generating summaries. We refer to the agent that uses our prompt as the ASL Agent.
Table 1: Comparison of different agents in terms of success rate in the AndroidWorld environment.
Type Agent Model Input Modality TSR $\uparrow$ Overall TSR $\uparrow$
Easy Medium Hard
Prompt-Driven SeeAct GPT-4o Image + Text 34.2 15.5 4.2 22.0
T3A GPT-4o Image 64.9 26.2 14.6 41.9
M3A GPT-4o Image + Text 60.5 20.2 8.3 36.6
ASL-Agent Qwen2.5-7B-Instruct Text 5.6 0.0 0.0 2.5
Fine-Tuned ASL-Agent Qwen2.5-7B-Instruct-SFT Text 72.2 26.8 3.13 42.5
Qwen2.5-7B-Instruct-ASL (Ours) Text 72.2 32.1 12.5 46.3
4.2 Online Comparison
In this experiment, we evaluated model performance in real-time interactive environments using the AndroidWorld rawles2024androidworld and AndroidLab xu2024androidlab datasets. This setup is well-suited for testing our core motivation: that models should go beyond syntactic patterns and learn the semantics of actions—that is, their actual outcomes. In online testing, instead of evaluating each step in isolation, it assesses success based on whether the user goal is reached. Specifically, AndroidWorld provides an emulator that mimics real-world conditions by automatically executing actions and employs a hand-crafted rule-based system to verify task completion by checking system status. For AndroidLab, we use 30 provided tasks, excluding those involving question answering. We evaluate AndroidLab by executing actions within the AndroidWorld emulator and using an LLM-based evaluator to determine whether each task is completed successfully.
Table 1 presents the results on AndroidWorld, reporting success rates across tasks of varying difficulty—categorised as easy, medium, and hard—as well as the overall success rate. As shown in the table, our method outperforms the baseline by a significant margin of 3.8% in overall success rate. The most substantial improvements are observed in medium and hard tasks. Notably, our model maintains strong performance across all difficulty levels, especially where reasoning and semantic understanding of actions are more challenging. We further evaluate our method on the AndroidLab benchmark, with results presented in Table 2. The results are consistent to the results on AndroidWorld. Specifically, our model outperforms the pure syntax learning baseline Qwen2.5-7B-Instruct-SFT, and achieves the highest success rate. Together, these results highlight the generalisation capability of our model in real-world mobile UI environments and underscore the effectiveness of our ASL framework.
Table 2: Comparison of different agents in terms of success rate in the AndroidLab environment.
Type Agent Model Input Modality TSR
Prompt-Driven AndroindLab-Agent Claude-3.5-Sonnet Image + Text 24.2
Gemini-1.5-Pro Image + Text 13.0
GPT-4o Image + Text 30.1
Gemini-1.5-Pro Text 17.2
GPT-4o Text 24.7
Fine-Tuned ASL-Agent Qwen-2.5-7B-Instruct-SFT Text 33.3
Qwen-2.5-7B-Instruct-ASL (Ours) Text 35.5
4.3 Offline Comparison
In this experiment, we evaluated model performance on the offline benchmark AndroidControl android_control . Specifically, we compared predicted actions against ground truth annotations. For action types, we assess whether the predicted action type string exactly matches the ground truth. For actions that involve coordinates, such as click and long press, we follow the evaluation protocol rawles2024androidinthewild to determine whether the ground truth coordinate falls within the bounding box of the predicted UI element.
AndroidControl provides four evaluation splits: IDD (In-Distribution Data), where all UI states and actions are sampled from the same distribution as the training data; Task-Unseen, where the task instructions are not observed during training; Cat-Unseen, where the UI categories are unseen; and App-Unseen, which involves entirely unseen applications. As shown in Table 3, we collect and evaluate both prompt-driven and models fine-tuned on AndroidControl.
Our model achieves the highest performance across all four splits. Under the IDD setting, it outperforms the strongest baseline with a significant margin. Moreover, it demonstrates strong generalisation to out-of-distribution scenarios. On the Task-Unseen split, our model surpasses the previous best. Similar improvements are observed on the Cat-Unseen and App-Unseen splits, highlighting the model’s robustness to new tasks, categories, and applications.
Table 3: Comparison across agents of step accuracy in the four splits of the AndroidControl test set.
Type Agent Model Input Modality SSR $\uparrow$
IDD Task-Unseen Cat-Unseen App-Unseen
Prompt-Driven SeeAct GPT-4o Image + Text 31.5 30.7 30.6 30.9
T3A GPT-4o Text 56.1 55.8 56.5 54.2
M3A GPT-4o Image + Text 60.8 59.3 60.8 60.4
ASL-Agent Qwen2.5-7B-Instruct Text 26.1 26.5 26.5 26.9
Fine-Tuned ASL-Agent Qwen2.5-7B-Instruct-SFT Text 67.4 61.2 59.7 59.7
Qwen2.5-7B-Instruct-ASL (Ours) Text 69.4 62.3 61.2 61.8
Table 4: Ablation study on the loss function.
$\ell^{SFT}_{\theta}$ $\lambda_{i}\,\ell^{SFT}_{\theta}$ $\ell^{ASL}_{\theta}$
OOD SSR (%) 49.8 49.9 51.7
Table 5: Ablation study on reward design.
LLM SEE (Ours)
OOD SSR (%) 50.6 51.7
4.4 Ablation Study
In this section, we study two key components of our ASL framework: the loss function and the reward model used during training. Ablation studies are conducted using the Qwen2.5-7B-Instruct model, trained on 3K samples from the AndroidControl training set and evaluated on 1K samples from the AndroidControl OOD test set. Results for the loss function ablation are presented in Table 5, where $\ell^{\text{SFT}}_{\theta}$ denotes the standard SFT loss used in the syntax learning paradigm, $\lambda{i},\ell^{\text{SFT}}_{\theta}$ represents the weighted SFT loss in our framework, and $\ell^{\text{ASL}}_{\theta}$ refers to the full loss function used in ASL. The results demonstrate that $\ell^{\text{ASL}}_{\theta}$ leads to the highest performance, highlighting its effectiveness in enabling the App agent to learn action semantics and improving robustness to OOD scenarios.
Table 5 reports the ablation results on the reward model. In this table, LLM denotes a baseline that directly uses Gemini-2.0-Flash to compute the semantics reward, while SEE refers to our proposed semantics estimator. The superior performance of SEE confirms its benefit in providing more informative and stable reward signals during training. These findings validate the effectiveness of both the loss function and the SEE reward model as core components of the ASL framework.
5 Conclusion
We present Action Semantics Learning (ASL), a novel framework for training App agents that prioritises semantic understanding of actions rather than syntactic reproduction. By defining action semantics as the UI state transition induced by an action, and introducing a lightweight semantic estimator (SEE) to provide reward signals during training, ASL enables agents to treat semantically equivalent actions equally, even when their textual forms differ.
We theoretically prove that ASL improves robustness against out-of-distribution (OOD) variations, and our extensive experiments confirm this. In online evaluation on AndroidWorld and AndroidLab, ASL-trained agents outperform both prompt-based GPT-4o agents and fine-tuned baselines, achieving higher task success rates, especially on medium and hard tasks that require flexible reasoning and semantic generalisation. On the offline AndroidControl benchmark, ASL significantly improves step-level accuracy under unseen task and app settings compared to syntax-based fine-tuning.
Limitation and future work. ASL currently relies on LLMs to simulate UI state transitions using language-based descriptions. A more concrete and detailed simulation that incorporates visual information would be a future focus.